RFBG Commissioning with Adaptive Secondary - 03 April 2010 UT

Night 15:

Observer: JHill, DMiller (Tucson)
Telescope Operator: JHill (Tucson)
Telescope Support: JUrban, MMidkiff
Adaptive Operator & Wind Monitor: RBriguglio (LBTO)
Vibration Operator: JEsquerra
W-Operator: EPinna (Arcetri)

Summary

We started the night with thin cirrus and medium seeing. Both clouds and seeing got worse as the night wore on. Nonetheless......

We completed an improved lookup table of Z6 versus elevation and implemented it on M2.

The first astronomical star was detected on the W-unit (in very poor seeing)!

We made some more tests of the new IRTC software,

Closed at 2AM for heavy clouds.

Preparations

JUrban changed out one of the HBS cooling pumps (#2).

Andrew reports: I'd like to try a combination of +15" M2 on-sky pointing tip and -5" M2 on-sky-pointing tilt.

Doug and Chris put in new versions of IRTC and irc software.

Details

Open

01:55 Open shutter doors - wind < 2 m/sec, T=0.8 degC, D=-22 degC, Thin cirrus clouds over most of sky

02:02 Open rear vent doors

02:03 Raise DX primary, but no air pressure. Jeff and Michael investigating. (perhaps it didn't get restarted after the power outage on Thursday morning) For the new compressor (running both mirrors) we only had to press the enable button on the ECSGUI,

02:10 Runa sets the shell at zenith

02:15 Home the DX M2 hexapod.

StartIRTC, but it gave an X Error on lbtdu16, so running it from my laptop. (I tried to change the date to 20100403 manually, but couldn't.)

02:31 pointto Pollux TRACK (1st try failed as rotator no ready)

02:36 pointto Pollux TRACK PAM- Adjust IE=-52 CA=-24 (with IA=-400) Coma was small at RX=-460 RY=+40.

02:44 takepic 0.01 irtc.20100403.024407.fits

02:45 takeseq 0.01 100 irtc.20100403.024438.fits cube 13

Now I was able to change the file prefix date on the IRTC GUI . (You can't change it while in TC mode.)

02:47 takeseq 0.01 100 irtc.20100403.024702.fits cube 14

02:50 pointto M5_0313 TelM=TRACK PAM- IA=-420 IE=-88 CA=-24 RX=-360 RY=-20

03:06 takeseq 1.0 100 irtc.20100403.030645.fits cube failed? = cube 15 arrived later

03:11 Applied +1 micron Z6, but was wrong sign

03:14 Increase to -1.5 micron Z6

03:15 -2.5 Z6 (forces won't allow us -3.0) This is summing with -2.5 already applied by lookup table (we think).

03:18 takeseq 1.0 100 irtc.20100403.031823.fits cube failed = cube 16 arrived later

03:25 move to zenith to turn off automatic Z6 and HO-correction

03:28 takeseq 1.0 100 (blank sky) irtc.20100403.032818.fits cube failed = cube 17 arrived later

03:30 takeseq 0.01 100 (blank sky) irtc.20100403.033037.fits cube 18

03:32 pointto M5_0313 TelM=TRACK PAM- IA=-430 IE=-95 CA=-24 image is quite elongated EL=19.8

03:33 -3.0 Z6 made it worse.

03:35 +3.0 Z6 was rejected, and shell RIP'd. (maybe was wrong mode)


03:48 Apply +5.0 Z6

03:49 Wind in dome (M2) increased to 3 m/sec. Closed Rear vent doors (all closed)

03:52 Apply +5.5 Z6 at 19.3 degrees = OK within seeing noise

03:53 remove the Z6 correction

03:54 pointto M5_0325 TelM=TRACK PAM- IE=-85

04:40 Apply +2.4 Z6 at EL=43.3 deg - seeing is worse so it is harder to evaluate

04:04 pointto M5_0347 TelM=TRACK PAM- IE=-83 CA=-28

04:05 start Right Mirror Ventilation and Left mirror ventilation

Runa measures 4 m/sec on balcony.

Adjust RX=-440 IE=-65

+1.2 Z6 for 58.8

These cubes have +1.4 Z6

04:24 takeseq 0.4 100 irtc.20100403.042435.fits

04:26 takeseq 1.0 100 irtc.20100403.042600.fits (succeed)

04:30 - 04:49 Doug took 13 images (20-32) with block and no-block, long and short exposures (0.001 to 1 second) and fast and slow frame rates (10000 to 5 Hz). They all returned success.

04:37 reset shell to see if edges improve - they were the same

04:49 Up to zenith for 5 minutes so Jorge can reset the vibration equipment

04:50 Runa sees 5 m/sec at M2, and 9 m/sec on patio. We must be pointing into the wind, so will choose targets father East.

04:55 John took another takeseq 1.0 100 which was also successful.

04:58 pointto M5_1332 TRACK PAM-

05:08 +2.0 EL=60.0 (somewhat better seeing than an hour ago, although still not especially good.)

05:10 pointto M5_0444 TRACK PAM- CA=-38 RY=-40 +1.5 Z6 at 66.3 deg

05:17 pointto M5_0441 TRACK PAM- +1.2 Z6 at EL=71.5

05:24 pointto M5_0425 TRACK PAM- RX=-500 RY=-60 IE=-53 CA=-39 0.0 +-0.3 at EL=85.8

05:34 pointto M5_1311 TRACK PAM- RX=-420 RY=-60 IE=-60 CA=-37 +2.0 at EL=61.1

05:44 pointto M5_0460 TRACK PAM- RX=-380 IE=-70 +3.0 at EL=53.2 deg

05:50 pointto M5_1371 TRACK PAM- RX=-380 RY=-20 IE=-90 CA=-25 +4.5 at EL=27.5

05:57 pointto M5_0387 TelM=TRACK PAM- IE=-80 1.7 Z6 at EL=55.8

06:03 pointto M5_0358 TelM=TRACK PAM-

06:05 pointto M5_1246 TelM=TRACK PAM- +2.0 at EL=54.0

06:13 pointto M5_0354 TelM=TRACK PAM- +3.6 at EL=37.5

06:18 pointto M5_0347 TelM=TRACK PAM- RX=-360 RY=-40 +3.2 at EL=42.8

This makes 14 (somewhat noisy points) in the Z6 versus Elevation plot. Much better than our previous 3 points.

06:xx pointto BS9143 gs=0 TelM=ACQUIRE

Interupt to practice closing and opening the dome

07:05 pointto BS9143 gs=0 TelM=ACQUIRE (again) - Acquire Movie is all black

07:10 stop GCSR; stopAGW -u 2; startAGW -u 2; home motors

07:15 pointto BS9143 gs=0 TelM=ACQUIRE (again) - OK guide camera is happy again after a power cycle. RX=-340 RY=-80 IE=-83 CA=-25

07:25 pointto BS9152 gs=0 TelM=ACQUIRE

07:28 Runa resets shell to apply the new lookup table values for Z6.

RX=-360 RY=-60 Z=2.7 CA=-33 IE=-74

07:42 pointto BS9152 gs=0 TelM=ACTIVE Pupil is very nice looking, no midscale aberrationsvisible to the eye. Seeing varying between 1 and 2 arcsec.

07:50 pointto BS9152 gs=0 TelM=ACQUIRE PAM-

Install Andrew's +15 Tip and -5 Tilt. This applies a correction to M2 of dX=+0.255 dY=-0.766 dZ=0 RX=147.3 RY=49.1 Unfortunately, this puts M2 beyond its motion limits. So we move M1 and M2 both +1 mm in Y. This leaves us very close to the M1 +Y limit at EL=70.

DX M1 Global X=-1.331 Y=+1.193 Z=0 RX=-11.6 RY=+2.5
DX M1 Total X=-0.392 Y=+2.846 Z=+0.087 RX=-7.47 RY=+21.19 at EL=70.5

DX M2 Global X=2.500 Y=-2.63 Z=2.8 RX=-360 RY=-60
DX M2 Total X=2.755 Y=-3.396 Z=2.800 RX=-212.7 RY=-10.9

08:00 pointto BS9152 gs=0 TelM=ACTIVE

08:11 pointto BS9152 gs=7 TelM=ACTIVE but now seeing is worse than 2 arcsec. Not useful for Andrew's field aberration measurements.

08:15 pointto BS9152 gs=12 TelM=ACTIVE

W-unit testing

08:18 pointto M5_1396 gs=2 ACTIVE - This is the first astronomical star detected on W-unit! We see live (and very turbulent) photons on both the technical viewer and the CCD39 pyramid camera. Seeing on the off-axis guider is bouncing between 1.5 and 2 arcsec.

Enrico says the default pointing (on-axis star position with off-axis guide star) was within 2 arcsec!

08:40 Guidestar lost to clouds

Close

08:50 Stop Guiding, Rest the Shell, Stop Primary Mirror Ventilation, Slew to EL=20.

08:57 Close Shutter Doors

09:02 Park telescope at AZ=90 EL=90. Stop HBS.

-- JohnHill - 03 Apr 2010
Topic revision: r5 - 04 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback